OpenStudy (anonymous):

Value of "n choose k" when 0

OpenStudy (dumbcow):

this is impossible....n must be greater or equal to k

OpenStudy (dumbcow):

if value must be given it would be 0

OpenStudy (anonymous):

And how you come to the conclusion that it is impossible? does the definition of n choose k say so?

OpenStudy (dumbcow):

not sure mathematically but think about it practically...if you only start with 3 things, then how will you choose a group of 5 from the 3 things?

OpenStudy (dumbcow):

http://en.wikipedia.org/wiki/Combination

OpenStudy (anonymous):

well, to my knowledge "n choose k" is not always equal to "choose k for n". That is only a sub part of using n choose k. IDK, maybe I'm wrong. But my textbook proves the existence of that, but I cant understand.

OpenStudy (anonymous):

My lecturers famous quote- "Never believe wikipedia definitions" :(

OpenStudy (dumbcow):

lol right thats why i said "not sure mathematically" this is higher math stuff, i usually dont spend too much time thinking about these things

OpenStudy (anonymous):

okay, Thank you very much for your help though

OpenStudy (anonymous):

if 0 < n < k, then (n choose k) = 0 a simple demonstration is, if you have 3 objects, how many ways can you select 5? well, the answer is 0. mathematically speaking (n choose k) = n! / [ (n-k)! k! ] since n < k, then n - k < 0 (n-k)! = (n-k) (n-k-1) (n-k-2) (n-k-3) .... loosely speaking (n-k)! is infinitely big. so [n! / (n - k)! k!] = (n!/k!) (1/(n-k)!) = 0 if n > 0, k < 0, then (n choose k) = 0 by the same argument the interesting case is when n < 0, and k > 0 then (n choose k) = n (n-1) (n-2) ... (n - (k - 1)) / k! example: (2 choose - 5) = (-2)(-3)(-4)(-5)(-6) / 5! = -6

OpenStudy (anonymous):

well, at least that would my thought anyways, though the case n <0, and k>0 shows up in binomial series representation

OpenStudy (anonymous):

consider Pascals triangle 1 1 1 1 2 1 1 3 3 1 1 4 6 4 1 each number is the sum of the 2 above it .... where do we get the 1s? 0 1 0 0 1 1 0 0 1 2 1 0 0 1 3 3 1 0 0 1 4 6 4 1 0 and where do we get the zeros? 0 0 1 0 0 0 0 1 1 0 0 0 0 1 2 1 0 0 0 0 1 3 3 1 0 0 0 0 1 4 6 4 1 0 0 so the solution with regards to Pascal is that 4 choose 7 is equal to 0, and also stuff like 4 choose -3 is 0 as well